English

A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P (not B)

Advertisements
Advertisements

Question

A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P (not B)

Sum
Advertisements

Solution

P(A) = 0.42, P(B) = 0.48

P(A and B) = P(A ∩ B) = 0.16

P(not B) = P(B’) = 1 – P(B)

= 1 – 0.48

= 0.52

shaalaa.com
Concept of Probability - Probability of 'Not', 'And' and 'Or' Events
  Is there an error in this question or solution?
Chapter 14: Probability - EXERCISE 14.2 [Page 307]

APPEARS IN

NCERT Mathematics [English] Class 11
Chapter 14 Probability
EXERCISE 14.2 | Q 17. (ii) | Page 307

RELATED QUESTIONS

If `2/11` is the probability of an event, what is the probability of the event ‘not A’.


In Class XI of a school 40% of the students study Mathematics and 30% study Biology. 10% of the class study both Mathematics and Biology. If a student is selected at random from the class, find the probability that he will be studying Mathematics or Biology.


In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing at least one of them is 0.95. What is the probability of passing both?


In a class of 60 students, 30 opted for NCC, 32 opted for NSS and 24 opted for both NCC and NSS. If one of these students is selected at random, find the probability that

  1. The student opted for NCC or NSS.
  2. The student has opted neither NCC nor NSS.
  3. The student has opted NSS but not NCC.

A die has two faces each with number ‘1’, three faces each with number ‘2’ and one face with number ‘3’. If die is rolled once, determine

  1. P(2)
  2. P(1 or 3)
  3. P(not 3)

A dice is thrown. Find the probability of getting a prime number


A dice is thrown. Find the probability of getting:

 2 or 4


In a simultaneous throw of a pair of dice, find the probability of getting:

8 as the sum


In a simultaneous throw of a pair of dice, find the probability of getting a doublet of prime numbers


In a simultaneous throw of a pair of dice, find the probability of getting a sum greater than 9


In a simultaneous throw of a pair of dice, find the probability of getting odd number on the first and 6 on the second


In a simultaneous throw of a pair of dice, find the probability of getting a total of 9 or 11


In a simultaneous throw of a pair of dice, find the probability of getting a total greater than 8.

 

In a single throw of three dice, find the probability of getting a total of 17 or 18.

 

Three coins are tossed together. Find the probability of getting exactly two heads


Three coins are tossed together. Find the probability of getting at least two heads


Two dice are thrown. Find the odds in favour of getting the sum  What are the odds against getting the sum 6?


What are the odds in favour of getting a spade if the card drawn from a well-shuffled deck of cards? What are the odds in favour of getting a king?

 

A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white .


A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is red or even numbered.


Fill in the blank in the table:

P (A) P (B) P (A ∩ B) P(A∪ B)
0.35 .... 0.25 0.6

There are three events A, B, C one of which must and only one can happen, the odds are 8 to 3 against A, 5 to 2 against B, find the odds against C


One of the two events must happen. Given that the chance of one is two-third of the other, find the odds in favour of the other.


A card is drawn at random from a well-shuffled deck of 52 cards. Find the probability of its being a spade or a king.


A card is drawn from a deck of 52 cards. Find the probability of getting an ace or a spade card.


In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing at least one of them is 0.95. What is the probability of passing both?


Three integers are chosen at random from the first 20 integers. The probability that their product is even is


Out of 30 consecutive integers, 2 are chosen at random. The probability that their sum is odd, is


A bag contains 5 black balls, 4 white balls and 3 red balls. If a ball is selected randomwise, the probability that it is black or red ball is


A box contains 6 nails and 10 nuts. Half of the nails and half of the nuts are rusted. If one item is chosen at random, the probability that it is rusted or is a nail is


Three numbers are chosen from 1 to 20. The probability that they are not consecutive is


In a certain lottery 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy two tickets.


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×